myininaya (myininaya):

\[\text{ The remainder of } \frac{1^{2013}+2^{2013}+\cdots + 2012^{2013}+2013^{2013}}{2014} \text{ is ... }\]

myininaya (myininaya):

So I only just began thinking about this question And this question comes from the almighty @mayankdevnani

OpenStudy (sparrow2):

2012 or 2013 in the degree?

OpenStudy (mayankdevnani):

2013

myininaya (myininaya):

\[\text{ The remainder of } \frac{1^{2013}+2^{2013}+\cdots + 2012^{2013}+2013^{2013}}{2014} \text{ is ... }\]

ganeshie8 (ganeshie8):

I'll just leave a small hint : Pair up

myininaya (myininaya):

\[\frac{1^{2013}+2013^{2013}}{2014}+\frac{2^{2013}+2012^{2013}}{2014} +\frac{3^{2013}+2011^{2013}}{2014}+ \cdots + \\ \frac{1006^{2013}+1008^{2013}}{2014} \\ +\frac{1007^{2013}}{2014}\]

ganeshie8 (ganeshie8):

All the terms except the last cancel anfd if there is any justice at all, 1007^2013 must be easy to reduce

myininaya (myininaya):

\[\text{ how do you show } \frac{a^{2013}+(2014-a)^{2013}}{2014} \text{ is 0}\]

myininaya (myininaya):

I mean the remainder

OpenStudy (sparrow2):